I am having trouble with simultaneous equations.
So how do you solve this simultaneous equation
[
2a-b+2c=-3
a+2b+c=7
3a-b-c=-8
]

Answer by TimothyLamb(4379)

---
2a - b + 2c = -3
a + 2b + c = 7
3a - b - c = -8
---
put the system of linear equations into standard form:
---
x = a
y = b
z = c
---
2x - y + 2z = -3
x + 2y + z = 7
3x - y - z = -8
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix-method:
x = a = -1.1
y = b = 3.4
z = c = 1.3
---
system matrix
2 -1 2
1 2 1
3 -1 -1
---
inverse of system matrix
0.05 0.15 0.25
-0.2 0.4 0
0.35 0.05 -0.25
---
determinant of system matrix = -20
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php